OpenStudy (anonymous):

-4x^5 ------- x^5-2x^6

OpenStudy (anonymous):

Find the GCF of the 2 terms in the denominator and take it outside the parentheses which is x^5: \[\frac{ -4x^{5} }{ x ^{5}(1-2x) }\] Simplify from here, what do u get?

OpenStudy (anonymous):

i dont know..

OpenStudy (anonymous):

What is x^5/x^5 ?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

Good, so you have: (-4x^5)/(x^5(1-2x)) - the x^5 term in the numerator cancels the x^5 term in the numerator, and leaves: \[\frac{ -4 }{ 1-2x }\]

OpenStudy (anonymous):

Do you understand how I got that?

OpenStudy (anonymous):

Do you know how to find the GCF of 2 terms?

OpenStudy (anonymous):

yes can you help with his one ? -x-x-42 -------- x-7

OpenStudy (anonymous):

-x^2-x-42 -------- x-7

OpenStudy (anonymous):

had it wrong sorry

OpenStudy (anonymous):

Do you know how to factor a trinomial: x^2-x-42?

OpenStudy (anonymous):

um no

OpenStudy (anonymous):

Ok, there are 2 ways to factor trinomials 1. Factoring 2. Using the Quadratic formula: \[x = \frac{ -b \pm \sqrt{b^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

Let us use factoring: Can you find 2 numbers that add together to give you -1 and multiply to give you 42?

OpenStudy (anonymous):

@chels226 are u there?

OpenStudy (anonymous):

Can you find 2 numbers that add together to give you -1 and multiply to give you -42?

OpenStudy (anonymous):

(x^2-x-42)/(x-7) = \[\frac{ (x-7)(x+6 }{ (x-7) })\] Like factor (x-7) in the numerator and denominator cancels - and gives you an Answer: (x-6). You are lacking the basic skills to attempt these questions I would suggest you review this topic in its entirety and get some help from your instructor.

OpenStudy (anonymous):

sorry i had to get off!